OpenStudy (anonymous):

5/6s+3/4=11/12 solve for s

OpenStudy (anonymous):

3/4 = 9/12 5/6s = 11/12 - 9/12 5/6s = 2/12 5/6s = 1/6 Multiply 6 on both sides. 5s = 1 s = 1/5